CHEMWORK What is the composition of a liquid for which the vapor phase above the liquid has a pressure of 355 torr at 25°C? Vapor Pressures at 25°C Pure pentane511 torr Pure hexane 150. torr Pentane mole fraction in the vapor Hexane mole fraction in the vapor- Submit Show Hints Submit Answer Try Another Version 8 item attempts remaining
Answer #1

Pentane mole fraction in vapor = 0.8174

Hexane mole fraction in vapor = 0.1826

Explanation

Let mole fraction of pentane in liquid = x

then, mole fraction of hexane in liquid = 1 - x

Total pressure = (mole fraction of pentane in liquid) * (vapor pressure of pure pentane) + (mole fraction of hexane in liquid) * (vapor pressure of pure hexane)

355 torr = (x) * (511) + (1 - x) * (150 torr)

355 torr = 511x + 150 torr - 150 x

355 torr - 150 torr = 511x - 150 x

205 torr = 361x

x = (205 / 361)

x = 0.568

mole fraction of pentane = x = 0.568

vapor pressure of pentane = (mole fraction of pentane in liquid) * (vapor pressure of pure pentane)

vapor pressure of pentane = (0.568) * (511 torr)

vapor pressure of pentane = 290.2 torr

Pentane mole fraction in vapor = (vapor pressure of pentane) / (total pressure)

Pentane mole fraction in vapor = (290.2 torr) / (355 torr)

Pentane mole fraction in vapor = 0.8174

Hexane mole fraction in vapor = 1 - Pentane mole fraction in vapor

Hexane mole fraction in vapor = 1 - 0.8174

Hexane mole fraction in vapor = 0.1826